A swimmers swims across a river 37.0m wide at 1.25m/s and swims back at 2.00m/s. what is the swimmers average speed?
Answer #1

Time taken by 1st shift

t 1 = 37 / 1.25 = 29.6 sec

Time taken by 2nd shift

t2 = 37 / 2 = 18.5 sec

Av speed

Vs = 37*2 / ( t1 + t2)

Vs = 1.54 m/s

Comment in case any doubt please rate my answer....